### v2

# Epistemic Bias Injection: Biasing LLMs via Selective Context Retrieval

Abstract
When answering user queries, LLMs often retrieve knowledge from external sources stored in retrieval-augmented generation (RAG) databases. These are often populated from unvetted sources, e.g. the open web, and can contain maliciously crafted data. This paper studies attacks that can manipulate the context retrieved by LLMs from such RAG databases. Prior work on such context manipulation primarily injects false or toxic content, which can often be detected by fact-checking or linguistic analysis. We reveal a more subtle threat, Epistemic Bias Injection (EBI), in which adversaries inject factually correct yet epistemically biased passages that systematically emphasize one side of a multi-viewpoint issue. Although linguistically coherent and truthful, such adversarial passages effectively crowd out alternative viewpoints and steer model outputs toward an attacker-chosen stance.
As a core contribution, we propose a novel characterization of the problem: We give a geometric metric that quantifies epistemic bias. This metric can be computed directly on embeddings of text passages retrieved by the LLM. Leveraging this metric, we construct EBI attacks and develop a lightweight prototype defense called BiasDef for them. We evaluate them both on a comprehensive benchmark constructed from public question answering datasets. Our results show that: (1) the proposed attack induces significant perspective shifts, effectively evading existing retrieval-based sanitization defenses, and (2) BiasDef substantially reduces adversarial retrieval and bias in LLM’s answers. Overall, this demonstrates the new threat as well as the ease of employing epistemic bias metrics for filtering in RAG-enabled LLMs.
1 Introduction
Retrieval-Augmented Generation (RAG) has become a de facto component of modern large language models (LLMs), enabling systems such as Llama-3 [1], GPT-4 [2], and DeepSeek-R1 [3] to incorporate external knowledge at inference time. By retrieving relevant passages from a corpus and injecting them into the model context, RAG improves performance on open-domain and specialized queries [4, 5, 6].
This paper is about a largely underexplored attack surface in RAG-enabled LLM systems which often rely on loosely controlled, web-scale sources rather than curated private knowledge bases. Since LLM outputs are shaped by the retrieved context, any manipulation of the retrieval results can directly influence the model’s responses.
Prior work has shown that adversaries can poison retrieved context of RAG by injecting malicious passages or prompts into the corpus [7, 8, 9, 10, 11], for example via edited Wikipedia pages, deceptive webpages, or news articles [12]. Such content (e.g., explicit falsehoods, structured prompt injections) within the malicious passages can induce attacker-desired outputs [7], trigger harmful tool usage [10], leak sensitive context [9], or even cause denial of service [8]. Prior defenses [13, 14, 15, 16, 17] therefore focus on detecting linguistic fingerprints or eliminating them in the retrieved passages.
We are concerned with a more subtle form of context manipulation: Even when all retrieved passages are factually correct or linguistically pleasant, an attacker can inject epistemic bias, i.e., LLM outputs systematically presents only one-sided perspectives. Analogous to selective reporting the fact in journalism, this manipulation skews the framing of the response without introducing any overtly malicious fingerprints. We formalize this threat as the Epistemic Bias Injection (EBI) attack (Fig. 1), in which adversaries inject epistemically biased yet truthful passages into the corpus so that they crowd out alternative viewpoints during retrieval. As a result, the LLM is steered toward a particular ideological or interpretative stance, undermining the objectivity and reliability of AI-augmented search and question-answering systems [12, 18, 19].
Figure 1: EBI attack in a typical RAG system. A corpus of passages is embedded and stored in a vector database. Given a query, the retriever returns the top-$k$ most relevant passages, which are then combined with the query and passed to the generator to produce the final output. The attacker injects adversarial passages into the corpus to corrupt the knowledge base, thereby influencing the contextual passages and the generator’s output. See Appendix D for more examples.
The key challenge posed by EBI is that it violates the assumptions underlying existing defenses. Unlike prior context manipulation attacks, EBI leaves no obvious fingerprints such as explicit instructions or verifiable false statements. Consequently, it evades known-answer detection [13, 14], fact-checking [15], model realignment [16], and even enhanced critical thinking [17] by the LLM itself. This makes EBI different from previously studied forms of RAG poisoning.
Our Approach: We give a novel geometric formulation to measure the level of epistemic bias contained in retrieved context and LLM outputs—specifically, a polarization score (PS). This score is computed by projecting text embedding vectors onto a "polarization" axis in the embedding space, which can be identified algorithmically.
The PS formulation lets us precisely define the vulnerability and create attacks. We show that using the PS value, attackers can generate adversarial passages that are simultaneously highly relevant to a target query, strongly biased in perspective, and difficult to distinguish from benign content. When injected into the knowledge base, these passages dominate retrieval results and systematically distort LLM responses. Evaluation on Llama-3, DeepSeek-R1-Distill-Qwen, and GPT-4.1 using open-source datasets [20, 21, 22] shows that existing perspective-aware sanitization defenses [23, 24, 25] are largely ineffective at blocking EBI attacks: they either retrieve a substantial fraction of adversarial passages, resulting in significantly biased answers, or aggressively filter content at the cost of discarding many useful benign passages.
We also show that PS formulation gives us a content sanitization defense. We develop BiasDef, a lightweight, plug-and-play method that augments relevance-based retrieval with PS-based statistical filtering. On the same benchmark, BiasDef-enabled RAG retrieves substantially more benign passages while reducing adversarial ones, yielding over a $6.2\times$ reduction in answer bias [^1] Bias is quantified as the shift in polarization score relative to the unattacked setting; see Sec. 2.2 for details.. It is important to note that BiasDef is not a universal defense against all forms of RAG posioning attacks. It can quantify epistemic bias of a passage in given a set of passages. The method can work with language models without retraining and needs no new datasets.
Contributions: Our contributions are as follows:
  • We introduce PS, a continuous and interpretable metric that quantifies epistemic bias in text embeddings.
  • We formalize EBI, a previously unaddressed context manipulation attack that biases LLM outputs using factually correct yet one-sided content.
  • We design BiasDef, a lightweight defense prototype that leverages PS to substantially reduce bias in LLM answers.
  • We construct a RAG benchmark spanning multiple LLMs and public datasets to evaluate both the vulnerability of retrieval algorithms and the effectiveness of BiasDef.
2 Problem Formulation
2.1 Primer on RAG
A typical RAG system is illustrated in Fig. 1. It has a retriever, a generator, and a vector database for knowledge storage. In this system, each passage $d$ in the corpus $D$ is first encoded into a fixed-dimensional embedding vector $E(d)$ using a pre-trained encoder $E(\cdot)$ (e.g., Sentence-BERT [26], GTR [27]). These embedding vectors are stored in the vector database.
When a query $q$ is issued, it is encoded into a query embedding $E(q)$ using the same or a compatible encoder. The retriever then performs a search between the query embedding and all stored passage embeddings to identify the top-$k$ vectors. Each retrieved vector is associated with an identifier or index, which allows the system to look up the corresponding original passage from the corpus or metadata storage. A typical retrieval algorithm aims to identify the $k$ most relevant passages with the highest similarity score (SS) to the query.
A typical SS is cosine similarity [^2] Cosine similarity for normalized vectors is directly proportional to their L2 distance. It is commonly used in vector databases [28]. between the embeddings of the query and passage. Formally, given a query $q$ and a passage $d$, it is defined as:
[…]
A higher SS indicates that the passage is more semantically aligned with the query, thereby increasing its likelihood of being included in the final contextual input to the LLM.
Once the top-$k$ relevant passages are retrieved, they are combined with the original query to form a structured input prompt for the generator. A common approach is to prepend the retrieved passages as contextual information, followed by the query posed as an explicit instruction [29]. Appendix A shows an example prompt in typical RAG process. This prompt is then fed into the LLM, which generates a grounded and contextually relevant answer. The LLM implicitly attends to the passages during decoding, and the quality of the output depends heavily on these retrieved content.
2.2 Quantifying Epistemic Bias
To be able to quantify bias in a presented viewpoint, one approach would be to involve collecting human annotation—for instance, ranking passage alignment from strongly opposing to strongly supporting. However, real-time and continuous human annotation for passages output by LLMs is expensive, if at all desirable. In this paper, we show that a geometric metric can be computed in an unsupervised way from a small set of anchor passages that have polarizing viewpoints.
In short, we seek a scalar metric that can distinguish passages expressing different viewpoints. Modern LLM sentence encoders map text into a fixed-dimensional embedding space, which encodes frequently occurring semantic and stylistic patterns. The whole passage also has an embedding vector which is influenced from its constituent sentences. While instance-specific details may be attenuated in the final embedding of the passage, latent attributes such as topic, sentiment, and stance consistent across many sentences remain in it.
Key Insight: Our hypothesis is that the difference in the viewpoint of a given set of text passages, which are otherwise similar (e.g. answering the same question), is geometrically discernibly in some direction of the embedding space. Identifying that direction and projecting embeddings onto it yields a scalar value that reflects both the orientation and intensity of a viewpoint, enabling quantitative comparison across passages.
To identify such a direction, we take a set of anchor passages with diverse viewpoints on the same topic. We obtain their embedding vectors by an LLM encoder and perform principal component analysis (PCA) [30] on them. The leading principal component—which geometrically maximizes the average variance of projections in the Euclidean $\ell_{2}$ sense—is identified as the polarization axis. In other words, this axis maximizes the separation among anchor embeddings.
Formally, given $J$ anchor passages $d_{\text{anchor},j}$ associated with a specific question, we compute their embeddings and stack them into a matrix $\mathbf{E}\in\mathbb{R}^{J\times\ell}$, where $e_{j}=E(d_{\text{anchor},j})\in\mathbb{R}^{\ell}$. The polarization axis $\mathbf{u}_{\text{polar}}$ is obtained by solving
$\mathbf{u}_{\text{polar}}=\mathop{\arg\max}_{\|\mathbf{u}\|=1}\mathrm{Var}(\mathbf{E}\cdot\mathbf{u}),$ (2)
which corresponds to the eigenvector associated with the largest eigenvalue of the sample covariance matrix of $\mathbf{E}$. This is the leading principal component, or the polarization axis.
Each new text passage can then be projected onto the polarization axis. For a passage $d$, its embedding projected onto this axis produces a scalar polarization score (PS):
$\mathrm{PS}(d)=\mathbf{u}_{\text{polar}}^{\top}E(d),$ (3)
Now, epistemic bias between the given passage $d$ and any reference passage $d_{\text{ref}}$ (e.g., a human-labeled neutral passage) is measured as the PS shift:
$\textbf{Epistemic Bias:}\quad\triangle\mathrm{PS}=\bigl|\mathrm{PS}(d)-\mathrm{PS}(d_{\text{ref}})\bigr|.$ (4)
This completes our proposed geometric formulation of the notion of bias. Previous work [31] informally stated that bias in information retrieval can be thought of as the distributional mismatch between the retrieved content and the ground truth distribution. In this paper, we concretized this abstract idea into a quantifiable divergence in PS values.
A useful property that the PS metric satisifies is local consistency. Consider two passages $d_{1}$ and $d_{2}$ about a topic and let $\{\mathbf{u}_{k}\}$ denote the orthonormal PCA basis. The squared $\ell_{2}$ distance between embeddings can be decomposed as:
$\|E(d_{1})-E(d_{2})\|_{2}^{2}$ $=\sum_{k}\bigl((E(d_{1})-E(d_{2}))^{\top}\mathbf{u}_{k}\bigr)^{2}$ (5)
$\geq\bigl(\mathrm{PS}(d_{1})-\mathrm{PS}(d_{2})\bigr)^{2}.$
Therefore, if two passages have high similarity—a small Euclidean distance in their embeddings, their difference in PS score is small. We call this property local consistency.
Figure 2: PS values versus 7-point Likert stance scores (annotated by GPT-5.2) for passages drawn from 50 randomly selected questions spanning social, cultural, and technological topics. On average, PS achieves a Spearman’s rank correlation [32] of 0.896 with Likert scores, demonstrating a strong monotonic alignment with judgments of viewpoint strength. This strong correlation indicates that PS values effectively capture the epistemic stance expressed in the passages.
Does PS score really capture bias? We performed the following experiment: We randomly selected a set of $50$ passages from a public dataset (details of which are given later) and then asked GPT-5.2 LLM to rank passages on how strongly they "support" or "oppose" the original question. We then compare it with our geometrically computed PS score, which is computed using embeddings from a base LLM different from the judge. Fig. 2 shows that the PS scores consistently correlate with the ranking given by the LLM acting independently as a judge. If one were to take GPT-5.2 as representative of how humans would rank passages, we can see that PS scores capture the notion of epistemic bias well.
This evaluation also shows that passages expressing supportive and opposing viewpoints tend to lie on opposite sides of the mean along the polarization axis, while neutral passages are near the mean. In this paper, we focus mainly on identifying binary viewpoint scenarios, capturing the most polarizing dimension (e.g., support vs. oppose), though we believe extensions to multiple axes are possible (see Sec. 7).
3 Threat Model and Attack Goals
In our threat model, the adversary’s sole capability is that it can inject a limited number of passages into the knowledge corpus used to populated the RAG system. Thus, it can poison some of the RAG database entries which can be domain-specific information used by a RAG-enabled LLM to answer a specific question. Taking any target open-ended question, we assume that the attacker aims to craft factually correct yet semantically biased passages that skew the retrieved context toward a particular stance. Once retrieved, these adversarial passages are incorporated into the LLM’s input context, which in turn can bias the generated answer output by the LLM.
Black-box access: We assume that the LLM encoder, the indexed benign passages, and the LLM model parameters are fixed and unavailable to the attacker; the adversary cannot modify the retriever, alter similarity metrics, or tamper with pre-existing entries in the database. Instead, the attack can only inject new adversarial passages. An adversarial passage $d_{\text{adv}}$ should have $3$ properties below to work effectively.
Property 1—High relevance: LLMs have a limited-length input context, especially in locally deployed small-scale settings (e.g., 4,000–32,000 tokens) due to constraints in model capacity, memory consumption, and generation latency. Hence, only a small subset of the retrieved passages from RAG database can be included in the input context. So, attackers must ensure their adversarial passages rank highly in their relevance to the query (i.e., SS values) during retrieval. Stated formally, we want: $\text{SS}(d_{\text{adv}})>\text{SS}(d_{\text{benign}})$, so that $d_{\text{adv}}$ is prioritized during retrieval over other benign passages $d_{\text{benign}}$.
Property 2—Epistemic bias: The passage must convey a specific ideological or opinionated stance that aligns with the attacker’s intended bias. This property is expressed as $\text{PS}(d_{\text{adv}})>(\text{or}<)~\text{PS}(d_{\text{benign}})$ for all benign passages $d_{\text{benign}}$. To maintain a consistent bias, multiple adversarial passages from a single attack must all have PS values either consistently higher or consistently lower than those of benign passages.
Property 3—Stealth (No explicit fingerprints): The adversarial passages must avoid introducing explicit or easily detectable fingerprints that distinguish them from benign content. In particular, they should remain factually correct, linguistically natural, and free of overt signals such as false statements, profanity, or instruction-like prompts. As a result, adversarial passages should blend into the distribution of benign passages under commonly used (non-PS) statistics.
4 EBI Attack on RAG Databases
Based on our metric formulation of PS, we propose an automated attack that produces adversarial passages for arbitrary query given. We leverage publicly available web sources and the generative capacity of LLMs to facilitate a fully automated attack workflow. It does not rely on human attackers.
Workflow of EBI attack: The core of EBI is to produce human-like yet epistemically biased text for a target question. Because the attacker controls the content fully, we can use an LLM to help generate these passages. Prior work suggests that dense retrievers tend to favor LLM-written content in retrieval [31, 33], so we can expect that using a state-of-the-art LLM to produce text will be effective. The attack process for a target question follows several steps:
  • (1) The attacker first needs $J$ passages with different perspectives on the question $q$, to serve as candidate passages for the attack. One approach is to retrieve these passages via keyword search using search engines. In practice, the attacker may prompt an LLM to generate them, using the Prompt-Synthetic in Appendix A as the prompt:
    $\{d_{\text{synt},j}\}_{j\in[1,J]}=\textbf{LLM}(\text{Prompt-Synthetic},q),$ (6)
    where LLM represents the LLM generation process. We use the latter.
  • (2) The attacker employs an encoder $E(\cdot)$ to project the candidate passages into the embedding space and performs PCA [30] on the embeddings $\{E(d_{\text{synt},j})\}_{j\in[1,J]}$, without requiring any viewpoint labels. In other words, these candidate passages also serve as the anchor passages to derive the polarization axis. By projecting the passage embeddings onto the resulting polarization axis, the attacker obtains their PS values $\{\text{PS}(d_{\text{synt},j})\}_{j\in[1,J]}$.
  • (3) The attacker selects the most biased passage $d_{\text{bias}}$ by choosing the passage with the highest or lowest PS:
    $d_{\text{bias}}=\mathop{\arg\max}_{d_{\text{synt},j}}\text{PS}(d_{\text{synt},j})\text{ or }\mathop{\arg\min}_{d_{\text{synt},j}}\text{PS}(d_{\text{synt},j})$ (7)
    aiming to induce a positive or negative shift in the generated answer along the polarization axis [^3] We note that the terms “positive” and “negative” are used solely to distinguish opposing viewpoints (e.g., supporting vs. opposing a policy), and do not imply any evaluative or emotional judgment about the content of the viewpoints themselves.. For convenience, we refer to these as positive EBI and negative EBI, respectively.
  • (4) The attacker then repeatedly queries the LLM to generate $n$ adversarial passages for the question $q$ that mimic the viewpoint of $d_{\text{bias}}$ while varying their presentation
    $d_{\text{adv},j}=\textbf{LLM}(\text{Prompt-Adv},q,d_{\text{bias}}),\quad j\in[1,n].$ (8)
    To ensure that the generated passages are factually accurate and semantically coherent, we carefully design the prompt (see Prompt-Adv in Appendix A) to guide the model toward grounded and contextually relevant content. We also control decoding parameters—such as using a low temperature and a high nucleus sampling threshold [34]—to reduce randomness in generation, thereby limiting unwarranted speculation and improving alignment with $d_{\text{bias}}$.
  • (5) Finally, the $n$ adversarial passages are injected into the database. Once retrieved, these passages introduce attack-desired epistemic bias into the generation process.
Satisfying attack conditions in practice: We explicitly design the attack workflow to satisfy each of the three attack conditions defined in Sec. 3. To satisfy,
  • Property 1 (High relevance), adversarial passages are generated in a query-conditioned manner, prompting the LLM to produce content that directly addresses the target question. This design encourages strong semantic alignment between the generated passages and the query in the embedding space [31, 33].
  • Property 2 (Epistemic bias), we guide generation using a biased passage selected from an extreme end of the polarization axis, thereby steering the LLM to consistently express a targeted viewpoint. Table 1 shows a property check result from our benchmark detailed in Sec. 5.2, which validates the effectiveness of our attack.
  • Property 3 (Stealth), we instruct the LLM to maintain factual correctness and linguistic coherence, avoiding explicit signals such as false statements or unnatural phrasing that could serve as detectable fingerprints.
5 Prior Sanitization Defenses vs. EBI Attack
5.1 Existing Sanitizers and Their Weaknesses
The most basic countermeasures against EBI to consider are those which rely solely on maximizing the relevance (i.e., SS value) of the retrieved passages. Along this line of defenses, dense-retrieval methods such as DPR [35] and Sentence-BERT [26], are widely adopted in RAG systems due to their effectiveness in capturing semantic similarity. However, they typically do not have any built-in defense mechanisms against EBI. Because retrieval decisions are driven solely by relevance, adversarial passages that are both biased and highly similar to the query can be easily retrieved. Our undefended baseline employs these mechanisms by default.
Property 1 & 2 Only Property 1 Only Property 2 Neither
Negative EBI 74.7% 16.2% 7.3% 1.8%
Positive EBI 78.0% 10.3% 8.7% 3.0%
Table 1: Among the 4,520 adversarial passages generated by our workflow for 452 queries in WIKI-BALANCE [20], more than 74% satisfy both Property 1 and Property 2.
In contrast to these relevance-only retrievers, several perspective-aware retrieval methods have been proposed to balance relevance, diversity, and conflict among the retrieved passages. They can serve as potential defense mechanisms for EBI attacks. Prominent examples of these include MMR [24] and SMART [25]. MMR reduces redundancy by penalizing candidates that are highly similar to already selected passages. It iteratively selects passages based on a relevance-diversity score, which accounts for both similarity to the query and dissimilarity to previously selected passages, thereby encouraging content diversity. SMART employs a more sophisticated selection strategy to optimize the trade-off among relevance, inter-passage similarity, and conflict resolution. Although these methods are effective in avoiding repetition and limiting the dominance of adversarial passages, as we will show later, they cannot fully eliminate EBI adversarial content—the diversity mechanism inadvertently encourages the retrieval of both benign and adversarial passages (see Fig. 3(a)).
Figure 3: Illustration of state-of-the-art retrieval strategies in the embedding space. (a) As indicated by the red arrows, MMR [24] and SMART [25] tend to favor diversity by selecting passages that are not necessarily the most similar to the query. (b) BRRA [23] retrieves passages relevant to both the original query and its noise-perturbed variants, then re-ranks them. As a result, benign passages near the perturbed queries (shown as green circles) may be retrieved even if they are distant from the original query.
Lastly, unlike the above methods, BRRA [23] introduces query perturbation to enhance diversity. It generates multiple noisy perturbations to the query and retrieves passages relevant to both the original and perturbed queries (illustrated as green circles in Fig. 3(b)). It then re-ranks all retrieved passages based on retrieval frequency and rank to promote coverage across the semantic space. BRRA is effective in low-dimensional embedding spaces, where noise-induced perturbations cause significant angular deviation. However, in high-dimensional settings, random noise induces only minor directional shifts after normalization, due to the concentration of measure on the unit hypersphere [36]. As a result, the similarity between original and perturbed queries remains high (e.g., >>0.9), yielding minimal variation in retrieved results.
5.2 Benchmark for Measuring EBI Attacks
We built an RAG benchmark to evaluate the effectiveness of existing methods against our proposed EBI attack.
Datasets: Our experiments are mainly based on an open-source dataset, WIKI-BALANCE [20], containing 452 open-ended questions and a corpus of 4662 real Wikipedia pages relevant to these questions. While using WIKI-BALANCE as the primary dataset, we also applied our attack and defense methods to Reddit-Dialogues [^4] Each query in this dataset includes a Reddit topic and previous responses of users, the RAG generates a new response based on retrieved passages from a given knowledge database. [22] and HotpotQA [^5] We use ChatGPT-5 to extend some close-form questions (e.g., ”Were Scott Derrickson and Ed Wood of the same nationality?”) to open-ended questions (”How did Scott Derrickson’s and Ed Wood’s cultural and national backgrounds shape their filmmaking styles and themes?”). [21] to show the generalization across multiple datasets.
[…]
  • Knowledge database: The database initially indexes benign passages from one of the three public datasets.
  • Retriever: We use msmarco-distilbert-base-tas-b [37] as the encoder for our dense retriever. This pre-trained model is fine-tuned on large-scale passage ranking datasets using contrastive learning and has demonstrated strong performance in dense retrieval tasks [26]. During retrieval, both queries and documents are encoded into dense vectors, and their relevance is computed via cosine similarity. The retriever first coarsely selects the top-$4\times k$ most relevant candidate passages from the knowledge database, and then applies existing methods to filter the final top-$k$ passages.
  • Generator: We deploy LLMs with various model sizes—including Meta-Llama-3-8B [38], DeepSeek-R1-Distill-Qwen-14B [39], and GPT-4.1 [2] (via the OpenAI API)—using the vLLM framework [40] to generate answers to open-ended questions. A prompt containing the query, the $k$ retrieved passages, and critical thinking instructions is constructed to initiate the answer generation process (see Prompt-Generation in Appendix A).
Attacker settings: As outlined in Sec. 4, the attacker deploys a DeepSeek-R1-Distill-Qwen-14B model to generate $J=8$ synthetic passages. These passages are then evaluated using a separate sentence-t5-xl [41] encoder to calculate their PS values, as the RAG’s own encoder is unknown to the attacker. The attacker selects two passages with the highest and lowest PS values as biased passages. Based on the biased passages, the attacker prompts the model to generate $n=10$ adversarial variants for positive EBI and another 10 for negative EBI and finally injects them into the knowledge database.
[…]
We set $k=5$ as the default number of retrieved passages throughout our evaluation. This choice follows common practice in RAG systems, where a small number of high-quality contexts is preferred to balance informativeness and computational efficiency. Specifically, typical retrieved passages are long articles (e.g, Wikipedia pages), many of which exceed 10K tokens. A larger $k$ would slow down inference, introduce more distractors (i.e., irrelevant content), and may even cause the total context length to exceed the input token limit of our locally deployed models.
Apart from the results about generalization across LLM models and datasets, we present attack efficacy results only for the most representative setting: DeepSeek-R1-Distill-Qwen-14B [39] as the generator and WIKI-BALANCE as the evaluation dataset [20]. This is a typical dataset featuring open and controversial questions across diverse topics. The complete code and data are released in our artifact.
Evaluation metrics: We use the following metrics:
  • Adversarial Recall(A-Recall)@k: The proportion of attacker-injected adversarial passages appearing among the top-$k$ retrieved results. This metric evaluates the effectiveness of the attack in getting adversarial content retrieved. A successful attack aims to maximize this rate to crowd out benign content and influence the generated answer toward the attacker’s desired perspective. Conversely, the defense mechanism aims to minimize it.
  • Recall@k: The proportion of correctly identified relevant benign passages among the top-$k$ results, assessed using Qrels [^6] Qrels (manual relevance judgments) are standard evaluation resources that contain query-document pairs labeled with binary or graded relevance scores. These labels are annotated by assessors following strict guidelines and are widely used as ground truth for evaluating retrieval quality.. A higher Recall@k implies that the retrieved context is more informative and aligned with the intended question.
  • Polarization Score (PS): As defined in Sec. 2.2, PS quantifies the viewpoint of a passage. PS shift—the magnitude of the difference between the PS values in the attacked and unattacked cases—measures the impact of attacks on the bias level of the content.
Compared baselines: We compare the potential baselines, including the defenses described in Sec. 5.1:
  • No defense (No Def.): Directly retrieves the top-5 passages with the highest cosine similarity to the query [42].
  • BRRA [23]: BRRA introduces Gaussian noise into the query embedding to generate multiple perturbed variants. we set the noise intensity to 1.0 (i.e., the perturbation magnitude equals the original embedding norm) to maximize variant diversity. Passages relevant to the original and perturbed queries are retrieved, and the union of these results is re-ranked based on retrieval frequency and rank across all queries. The top-5 passages after re-ranking are then selected as the contextual input to the LLM.
  • MMR [24]: We deploy the native MMR algorithm and set $\lambda$—the coefficient that balances the relevance and diversity of the retrieved passage—to a commonly used value: 0.5.
  • SMART [25]: As described in [25], we construct a conflict-aware kernel matrix based on three components: (i) the query-context relevance matrix (measured via cosine similarity between each passage and the query), (ii) the similarity matrix between passages, and (iii) a conflict matrix derived NLI predictions. For conflict detection, we use a pre-trained NLI model [^7] Available at https://huggingface.co/MoritzLaurer/mDeBERTa-v3-base-xnli-multilingual-nli-2mil7.. This kernel matrix is then used to the described algorithm for passage ranking.
  • Automated Fact-Checking [15]: Existing automated fact-checking APIs provide per-claim (i.e., a complete sentence) check yet; thus, they cannot apply to long adversarial passages directly. We empirically compare the benign and adversarial passages, broken down into sentences, to test if fact-checking can distinguish between the two.
5.3 Retrieval Performance under EBI Attacks
Method Unattacked #Negative Injection #Positive Injection
1 5 10 1 5 10
No Def. 0 0.19 0.93 0.98 0.20 0.92 0.97
MMR 0 0.19 0.22 0.27 0.19 0.23 0.29
SMART 0 0.19 0.21 0.23 0.18 0.21 0.23
BRRA 0 0.04 0.19 0.20 0.04 0.19 0.20
Table 2: Average A-Recall@5, the proportion of adversarial passages in top-5 retrieved passages, for varying number of attacker-injected passages.
We evaluate how EBI attacks affect different retrievers by examining two key aspects on the top-5 passages retrieved by the retrievers: A-Recall@5 and Recall@5. An effective defense should minimize A-Recall@5 as much as possible, without sacrificing the retrieval quality measured by Recall@5.
A-Recall@5: Table 2 presents the average A-Recall@5 across all 452 WIKI-BALANCE queries under different injection intensities, where “#Positive Injection” and “#Negative Injection” denote the numbers of injected adversarial passages for positive EBI and negative EBI respectively. As the injection intensity increases, all retrievers exhibit higher A-Recall@5. Among the baselines, the No Def. baseline demonstrates the highest vulnerability to EBI attacks, attributable to its exclusive reliance on query-passage similarity for ranking. In contrast, MMR, SMART, and BRRA incorporate diversity criteria (Sec. 5.1) that partially mitigate adversarial retrieval, yet still leave over $19\%$ of adversarial passages unsanitized at injection intensities 5 and 10. This is nearly a $5\times$ reduction compared to No Def., but it comes at the cost of reduced Recall@5 (see Table 3).
Method Unattacked #Negative Injection #Positive Injection
1 5 10 1 5 10
[…]
Table 3: Average Recall@5, the proportion of relevant passages (as labeled by Qrels) among the top-5 retrieved passages. The $\pm$ values for "Unattacked" indicate the 99% confidence interval, quantifying the sampling noise.
[…]
Worst-case performance: Since the attacker can vary the number of injected adversarial passages, we compute the maximum ratio between A-Recall@5 and Recall@5 across all injection intensities to assess the worst-case performance of each defense method. Specifically, the maximum ratios are as follows: 98 (at #Neg=10) for No Def., 0.85 (at #Neg=1,5 and #Pos=5,10) for MMR, 0.86 (at #Neg=1) for SMART, and 20 (at #Neg=10) for BRRA. These results demonstrate that both No Def. and BRRA are highly susceptible to adversarial passage domination, with adversarial content frequently overwhelming the top-5 retrievals. In contrast, MMR and SMART reduce this vulnerability; however, they still allow a number of unsanitized adversarial passages that is comparable to the number of relevant benign passages.
Method Unattacked Avg. ||PS|| #Negative Injection #Positive Injection
1 5 10 1 5 10
[…]
Table 4: Average PS shift of the top-5 retrieved passages, expressed as a percentage of the unattacked Avg. $|\text{PS}|$.
Method Unattacked Avg. ||PS|| #Negative Injection #Positive Injection
1 5 10 1 5 10
[…]
Table 5: Average PS shift of answers generated by DeepSeek-R1-Distill-Qwen-14B [39], expressed as a percentage of unattacked average $|PS|$. The answers show a significant PS shift due to the biased context passages retrieved.
Figure 4: Both answer and retrieval PS values demonstrate strong positive correlation, with answer PS consistently mirroring viewpoint trends in retrieved passage PS.
In summary, existing methods either retrieve a high proportion of adversarial passages or sacrifice on useful benign passages in an effort to reduce adversarial ones.
5.4 Impact on the LLM Answers
Unattacked average absolute PS (Avg. ||PS||): This value, as reported in Tables 4 and 5, measures the original bias level in unattacked cases. It also indicates whether the defender misses benign passages with strong perspectives, resulting in more neutral but ambiguous content. BRRA, MMR, and SMART yield lower Avg. ||PS|| than No Def. for both retrieved passages and answers. This is because the diversity mechanism may aggressively attenuate strong perspectives.
PS shift caused by the attack: The attacker intends to make the PS of the retrieved passages biased toward its desired direction (i.e., ascending direction for positive EBI and descending direction for negative EBI). We report the average PS shift (see Eq. (4)) of the top-5 retrieved passages. The EBI attack causes the retrievers to retrieve biased passages, and their average PS shift increases as more adversarial passages are injected (see Table 4). This demonstrates the vulnerability of retrieval systems to deliberate EBI attacks through adversarial passages.
We also report on the average PS shift of the generated answers to every question in the corpus. As listed in Table 5, the EBI attack successfully induces PS shifts in the answers, and the magnitude of the shift increases with the number of injected adversarial passages. Moreover, the generated answers are positively correlated with the retrieved passages in terms of PS values. Fig. 4 quantifies such correlations.
Retrieval bias propagates through the generation process. LLM models fail to mitigate the injected biases during answer synthesis, even when equipped with critical thinking instructions in the prompt (see Prompt-Generation in Appendix A for the prompt used).
Vulnerability across models and datasets: A natural question arises: Does the impact of EBI remain if we switch to a different—or larger—LLM models? To answer this, we evaluate multiple LLMs using identical retrieval contexts, thereby isolating the model’s role in absorbing contextual bias. Table 6 presents representative results for the undefended retriever on WIKI-BALANCE.
Our results show that all three LLMs exhibit comparable PS shifts in their generated answers when under attack. The results in Table 6 also demonstrate that our attack successfully induces PS shifts in question answering across diverse datasets. As the number of injected adversarial passages increases, the answers exhibit stronger bias, reflected by larger PS shifts across all the datasets.
Pre-trained LLMs, regardless of size or architecture, remain similarly vulnerable to EBI attacks. Merely scaling up the model or switching architectures does not inherently improve robustness.
5.5 Ineffectiveness of Fact-Checking
To investigate whether existing automated fact-checking is a potential defense for EBI attacks, we compare both benign and adversarial passages to see if they are distinguishable with a standard fact-checking API. Specifically, we run the WikiCheck [15] API which return a result of “SUPPORTS” (TRUE), “REFUTES” (FALSE), or “NOT ENOUGH INFO” for each sentence in a passage.
As shown in Fig. 5(a), benign and adversarial passages give a comparable proportion of sentences labeled as “NOT ENOUGH INFO” by the fact-checking API. This observation can be attributed to two main factors. First, a fraction of sentences are not genuine claims but rather contextual or descriptive statements—such as definitions, methodological descriptions, or references—that do not convey verifiable factual assertions. Second, even when a sentence constitutes a claim, its validity often hinges on specific contextual scopes (e.g., internal experiment settings, unpublished data, or domain-restricted facts), while the general-purpose knowledge base leveraged by the fact-checking API lacks the corresponding evidence required for verification.
Verifiable claims in benign and adversarial passages are labeled as “SUPPORTS” or “REFUTES” by the API. Fig. 5(b) shows the distribution of support ratio (i.e., SUPPORTS / (SUPPORTS + REFUTES)). We see that adversarial passages exhibit a higher proportion of “SUPPORTS” labels. This phenomenon shows that our adversarial passages generated by LLMs use evidence that fact-checkers deem as facts.
Our attack creates passages that are as “factual" or more than benign passages to the WikiCheck fact-checker.
As a noteworthy caveat, all passages, including benign ones, contain a non-negligible fraction of “REFUTES” claims. This can arise from at least two factors. First, certain claims are only valid within the local context of the passage; when isolated and checked independently, the per-sentence fact-checker may misclassify them as false statements. Second, this can be because the current fact-checking API is not perfectly reliable and has occasional false negatives.
Method Unattacked Avg. ||PS|| #Negative Injection #Positive Injection
1 5 10 1 5 10
[…]
Table 6: Average PS shift across different models and datasets. The proposed EBI attack successfully biases the answers of prevalent LLM models with various parameter sizes across different datasets.
Figure 5: Fact-checking results: (a) Adversarial passages contain comparable proportions of sentences labeled as “NOT ENOUGH INFO” (0.69 and 0.68 on average for positive and negative EBI attacks, respectively), which are similar to that of benign passages (0.71). (b) Among verifiable claims, adversarial passages contain higher proportions of sentences labeled as “SUPPORTS” (0.73 and 0.79 on average for positive and negative EBI attacks, respectively) than benign passages (0.63).
6 BiasDef: A Principled, Adaptive Defense
Figure 6: Workflow of BiasDef. (a) The $S$ value space is searched to identify the Max-KL boundary $t_{ss}^{*}$ that maximizes the KL divergence between the PS distributions of the two subsets partitioned by $t_{ss}^{*}$. (b) We refine the passage set $\mathcal{S}(t_{ss}^{*})$ by removing false positives with outlier PS values and (c) recover false negatives through Mahalanobis-distance-based classification.
We have seen that existing sanitizers can attenuate the impact of EBI, but remain fundamentally insufficient. Is there an adaptive defense that can defeat EBI attacks as formulated?
To explore this question, we design of a more effective defense that explicitly leverages PS-based viewpoint quantification. The rationale behind our design is as follows: Although attackers leave no explicit surface-level fingerprints, adversarial passages exhibit statistically different distributions from benign passages when observed jointly in the PS-SS space.
To test this hypothesis, we design a defense prototype, BiasDef, which relies on explainable statistical methods to filter out adversarial passages. In this process, we do not assume that adversarial and benign passages are perfectly separable; instead, we explicitly consider the more realistic case where their distributions partially overlap. The reasons are as follows. First, as shown in Table 1, while many adversarial passages exhibit extreme PS values and higher SS values than benign passages, a non-negligible fraction does not. This violates the assumption that statistical distributions of adversarial and benign passages are perfect separable. Second, note that the defender may not know the exact anchor passages used by the attacker to construct the attack. So, our proposed BiasDef derives its polarization axis from a different set of anchor passages retrieved from the corpus, which naturally induces a PS distribution that differs from that used by the attacker.
Therefore, BiasDef is composed of two components: (1) a base detector that admits a theoretical performance guarantee in the idealized case where adversarial passages are perfectly separable from benign ones (i.e., all adversarial passages satisfy Properties 1 and 2); (2) a set of statistical countermeasures that mitigate false positives and false negatives of the base detector in realistic settings where adversarial and benign distributions overlap, i.e., they are not perfectly separable.
6.1 Base Detector
BiasDef identifies and filters adversarial passages by considering the PS score within a constrained similarity range. Given a query, we first use the default retriever to obtain an expanded candidate set that contains more passages than the final top-$k$ context budget. For these candidate passages, we jointly characterize each passage by its SS to the query and its PS value, forming a two-dimensional PS-SS feature space. Specifically, BiasDef computes a polarization axis by performing PCA on the embeddings of the retrieved candidate passages, and projects each passage onto this axis to obtain its PS value. This joint representation enables BiasDef to reason about epistemic bias conditioned on relevance, rather than treating relevance and bias as independent signals. To facilitate detection, we introduce the following definition.
Definition: Given a set $D_{\text{benign}}$ of benign passages and a set $D_{\text{adv}}$ of adversarial passages, we define $\mathcal{S}=\mathcal{S}(t_{ss})$ as a subset of the entire set $D_{\text{benign}}\cup D_{\text{adv}}$ that includes all passages whose SS exceeds a threshold $t_{ss}$. The gray area in Fig. 6(a) visualizes an example of such a subset. We denote the complement of this subset as $\overline{\mathcal{S}}=(D_{\text{benign}}\cup D_{\text{adv}})\setminus\mathcal{S}(t_{ss})$. Let $P_{\mathcal{S}}$ and $Q_{\mathcal{S}}$ denote the PS distributions of $\mathcal{S}$ and $\overline{\mathcal{S}}$, respectively. To facilitate computation, we discretize the PS distributions of both $\mathcal{S}$ and $\overline{\mathcal{S}}$. Let the PS axis be partitioned into $m$ equal-width bins [^8] We assume $m$ is sufficiently large (i.e., the bins are narrow enough) such that no bin contains both adversarial and benign passages when their PS values fall into disjoint intervals, as described in Property 2.. Define $c_{\mathcal{S}}(i)$ as the number of passages in $\mathcal{S}$ that fall into bin $i$, and similarly $c_{\overline{\mathcal{S}}}(i)$ for $\overline{\mathcal{S}}$. Then, the discrete PS distributions [^9] We add a small constant $\epsilon$ in piratical cases to avoid $\log(0)$ and $\log(\infty)$ when computing the KL divergence. See Algorithm 1 for details. are computed as:
${P}_{\mathcal{S}}(i)=c_{\mathcal{S}}(i)/\sum_{i}c_{\mathcal{S}}(i),~{Q}_{\mathcal{S}}(i)=c_{\overline{\mathcal{S}}}(i)/\sum_{i}c_{\overline{\mathcal{S}}}(i).$ (9)
The Kullback–Leibler (KL) divergence [43] between these two distributions is as follows:
$f_{\text{KL}}(t_{ss})=KL[P_{\mathcal{S}}\|{Q}_{\mathcal{S}}]=\sum_{i=1}^{m}{P}_{\mathcal{S}}(i)\log({\frac{P_{\mathcal{S}}(i)}{Q_{\mathcal{S}}(i)}}).$ (10)
Based on these definitions, the following theorem that defines the security property our base BiasDefdetector achieves.
Theorem 1. Suppose two conditions hold: (1) All passages in $D_{\text{adv}}$ have higher SS values than any passage in $D_{\text{benign}}$ (Property 1); and (2) All passages in $D_{\text{adv}}$ have PS values that are either greater than or less than those of all passages in $D_{\text{benign}}$ (Property 2). Then, the KL divergence $f_{\text{KL}}(t_{ss})$ in Eq. (10) achieves a local maximum at any threshold $t^{*}_{ss}\in(0,1)$ such that: $\mathcal{S}=D_{\text{adv}}$ and $\overline{\mathcal{S}}=D_{\text{benign}}$.
Proof.
A detailed proof is provided in Appendix B. ∎
This theorem gives us a defense procedure if the adversarial and benign distributions are perfectly separable.
Theorem-guided base detector: Theorem 1 says that any threshold $t^{*}_{ss}$ that cleanly separates $D_{\text{adv}}$ from $D_{\text{benign}}$ is guaranteed to be a local maximizer of the KL divergence $f_{\text{KL}}(t_{ss})$. Guided by this insight, our defense searches over the entire SS threshold space $t_{ss}\in[0,1]$ monotonically to identify the point where $f_{\text{KL}}(t_{ss})$ reaches a maximum away from the boundaries (i.e., excluding 0, 11):
$\textbf{Max-KL Boundary:}\quad t^{*}_{ss}=\arg\max_{t_{ss}}f_{\text{KL}}(t_{ss}),$ (11)
$s.t.\quad f_{KL}(t^{*}_{ss})\geq f_{KL}(t_{ss}),\;t_{ss}\in[t^{*}_{ss}-\delta,t^{*}_{ss}+\delta]$
Note that $f_{\text{KL}}(t_{ss})$ may be non-smooth in practice, particularly under finite-sample settings, and thus may have multiple local maxima. We select the first largest maxima found by decreasing $t_{ss}$ from $1$ down to 0. We empirically find that this often yields a separation boundary which aligns well with the ground truth boundary between adversarially crafted passages and benign passages, making it effective.
6.2 Mitigating False Positives and Negatives
As explained earlier, adversarial and benign passages may not be perfectly separable, i.e, some passage may be deemed both adversarial and benign with non-negligible probability. In such cases, the base detector—which is optimal in the perfect separable case—will have false positives or negatives. The following mitigation augments the base detector to deal with the possibility of false positives and negatives.
Mitigation of false positives: To mitigate false positives—benign passages with SS values comparable to those of adversarial passages (see Fig. 6(b))—we identify a subset $\alpha\subset\mathcal{S}=\mathcal{S}(t^{*}_{ss})$ consisting of passages whose PS values are closest to those in $\overline{\mathcal{S}}$ [^10] If the average PS of $\mathcal{S}$ is higher than that of $\overline{\mathcal{S}}$, we select passages in $\mathcal{S}$ with the lowest PS values; otherwise, we select those with the highest.. We then compare the KL divergence before ($KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]$) and after ($KL[P_{\mathcal{S}\setminus\alpha}\|Q_{\mathcal{S}\setminus\alpha}]$) excluding $\alpha$. If removing $\alpha$ increases the KL divergence, i.e., $\mathcal{S}\setminus\alpha$ and $\overline{\mathcal{S}}\cup\alpha$ are more statistically divergent, it suggests that $\alpha$ contains passages that are more like benign passages. We iteratively expand $\alpha$ until $KL[P_{\mathcal{S}\setminus\alpha}\|Q_{\mathcal{S}\setminus\alpha}]$ reaches a local maximum and begins to decrease. The final $\alpha$ is treated as a group of false positives and removed from $\mathcal{S}(t^{*}_{ss})$.
Mitigation of false negatives: To recover false negatives—those adversarial passages with low SS values—we apply a Mahalanobis-distance-based classification [44], a classical technique widely used for identifying outlier in high-dimensional embedding spaces [45]. This method captures the covariance structure of the detected adversarial passages in $\mathcal{S}(t^{*}_{ss})$ and identifies samples that exhibit small deviations from the cluster center. Specifically, a passage $d$ is considered a false negative and is included in the refined subset if its embedding $E(d)$ satisfies the following condition:
$\sqrt{(E(d)-\mu)V^{-1}(E(d)-\mu)^{\top}}<T$ (12)
where $\mu$ and $V$ denote the mean vector and covariance matrix of the embeddings of passages in $\mathcal{S}(t^{*}_{ss})$, respectively. This condition evaluates the Mahalanobis distance between $d$ and the adversarial cluster. The threshold $T$ is selected via hyperparameter tuning based on validation performance.
We adopt Mahalanobis distance because its squared form equals a weighted [^11] The unweighted version corresponds to the square of the standard $\ell_{2}$ distance. The weights adjust for different variation along different directions in the vector space, as in the sample covariance matrix $V$. sum of squared differences across embedding dimensions. In our setting, consequential false negatives are those which induce bias, and thus will have PS value that devates from zero. Therefore, they lie near the adversarial cluster along the principal component identified by PCA. Mahalanobis distance captures this: Eq. (12) checks the PS deviation of $d$ relative to the adversarial cluster.
After excluding the suspicious adversarial passages—those identified by first removing the false positives in $\alpha$ from $\mathcal{S}(t^{*}_{ss})$ and then adding the false negatives that satisfy Eq. (12)—from the retrieved candidate set, we select and return the final top-$k$ relevant passages. The defense pseudo-code is summarized in in Appendix C.
6.3 Effectiveness Evaluation of BiasDef
We now evaluate the efficacy of BiasDef against our proposed attack on the same benchmark described in Sec. 5.2.
A-Recall@5: BiasDef achieves an average A-Recall@5 of 0, 0.04, 0.19, 0.20, 0.04, 0.19, 0.20 for the unattacked case, #Negative Injection = 1, 5, 10, and #Positive Injection = 1, 5, 10, respectively, which are comparable to the lowest values achieved by the baselines (BRRA in Table 2). As shown in Fig. 7, BiasDef is able to completely filter out adversarial passages from the top-5 results for over 70% of the queries.
We further analyze the evasion rate of adversarial passages—the proportion of injected adversarial passages that successfully bypass BiasDef’s filtering mechanism. As shown in Table 7, BiasDef is less effective at filtering out those adversarial passages that do not simultaneously satisfy both Property 1 and Property 2. This is theoretically expected too, as such atypical passages do not produce significant KL divergence changes during the defense process.
Figure 7: CDF of BiasDef’s A-Recall@5. BiasDef completely filters out adversarial passages for over 70% queries.
Passage Properties #Negative Injection #Positive Injection
1 5 10 1 5 10
Property 1 & 2 0.29 0.20 0.11 0.26 0.19 0.09
Other 0.86 0.60 0.31 0.93 0.67 0.32
Table 7: Evading rate of adversarial passages under BiasDef’s defense. Adversarial passages that lack either Property 1 or 2 are more likely to evade the detection.
[…]
Unattacked average absolute PS (Avg. ||PS||): BiasDef does not significantly reduce the unattacked Avg. ||PS|| (0.069) compared with No Def. (0.060), indicating that BiasDef does not falsely filter out more benign passages with strong perspectives and high PS values.
[…]
Unlike alternative methods such as [16] and [46], BiasDef functions as a purely plug-and-play filter, requiring no modifications to the underlying LLM. Moreover, BiasDef introduces a modest additional cost (avg. of 465 ms) for mitigating adversarial bias. This overhead is negligible compared to the typical generation time of LLMs, which is several seconds in our setup. This design makes it lightweight, easy to deploy, and compatible with diverse RAG systems.
7 Limitations, Discussion, and Future Work
LLM vs. Human Judgement: Our formulation provides a quantitative measure of epistemic bias. We empirically see that passages with substantially different PS values tend to express different viewpoints as deemed by an LLM acting as a judge. As per the subjective and qualitative assessment by us (authors), the PS value does correlate with "support" vs. "oppose" viewpoint and strongly biased passages have high PS value magnitudes. This observation is consistent across datasets and queries considered in our experiments.
An important open question is how PS correlate with human assessment, and how does that affect human trust, perception, and belief formation. In particular, it remains unclear what magnitude of PS shift is sufficient to induce a meaningful change in a user’s interpretation or confidence in the generated answer. A comprehensive evaluation of these downstream human effects requires controlled user studies and is beyond the scope of this paper, but represents an important direction for future work.
Multi-dimensional extensions: Real debates can span multiple, sometimes orthogonal axes, beyond that considered in our PS. We used the most polarizing dimension, captured by the first principal component during PCA, which typically represents the strongest binary disagreement (e.g., support vs. oppose). Nonetheless, the approach can be extended to multiple dimensions by considering additional principal components. In principle, BiasDef could also be applied iteratively along these dimensions [47] to mitigate multi-axis polarization, but exact details remain future work.
General-purpose defense: RAG systems operate in complex environments that may be subject to multiple types of attacks, of which EBI is only one. A promising future direction is to integrate complementary defenses (e.g., BiasDef and KAD) into a unified, general-purpose defense system capable of mitigating a broader range of threats.
Method Unattacked Avg. ||PS|| #Negative Injection #Positive Injection
1 5 10 1 5 10
[…]
8 Related Work
8.1 Context Manipulation with Fingerprints
Context manipulation attacks aim to corrupt the retriever’s results by adding malicious content to the corpus. PoisonedRAG [7] and BADRAG [8] generate malicious passages that are highly relevant to a trigger question, using LLM prompting and contrastive learning, respectively. While these approaches increase the retrieval relevance of adversarial passages, their malicious content and generation pipelines lack semantic-level refinement, which can result in unnatural expressions or factual inaccuracies.
Beyond injecting malicious content, prompt injection attacks [48, 10] aim to trigger more harmful actions by inserting malicious instructions. These prompts override the system’s original task and coerce the LLM into executing an attacker-injected workflow. A recent adaptive attack, DataFlip [49] bypasses KAD [13, 14] defenses by specifically identifying the presence of its detection instructions.
Unlike the above approaches on constructing injected texts, AgentPoison [11] specially craft optimal “trigger” such that whenever a user’s query contains a secret phrase, the retriever will tend to retrieve the attacker’s documents.
Overall, these studies show that simply poisoning the corpus can mislead a RAG’s outputs. However, prior attacks typically rely on overtly malicious content (e.g., factually incorrect statements [7, 8], spam-like text with triggers [11, 50], or structured prompts [48]), which risks detection [13, 14, 16, 15, 17].
8.2 Other Forms of Bias Injection Attacks
Recent work [23, 51, 52] has studied other forms of bias injection into LLM outputs. These studies primarily focus on discriminatory (e.g., gender or racial) bias and conceptual bias (e.g., covert brand-name replacement for advertisement), rather than epistemic bias. More importantly, their threat models differ substantially from ours. Existing approaches either assume a stronger attacker capability [23]—which is impractical for external attackers—or operate at the training stage [51, 52], whereas our attack targets the retrieval stage. Specifically, [23] assumes that the RAG system’s internal structured prompt between the retriever and the LLM—used to instruct the LLM to generate responses based on retrieved context—can be directly manipulated. It further directly perturbs adversarial passage embeddings, which may introduce semantic errors or factual inconsistencies after decoding. [51, 52] focus on corrupting training data so that trained models learn biased knowledge.
8.3 Purely Retrieval-Based Defenses
Some work [53] has shown that retrievers relying solely on semantic relevance—such as BM25 [54], DPR [35] and Sentence BERT [26]—may return inherently biased documents, limiting the diversity and usefulness of the retrieved results. Such bias in retrieval can lead to the omission of helpful counter-perspectives or even skew the LLM’s output.
To improve perspective awareness, several studies have proposed perspective-aware retrievers (e.g., MMR [24], BRRA [23], and SMART [25]). We provide a detailed description of these methods in Sec. 5. While these methods are effective at avoiding repetition and reducing the dominance of adversarial passages in the top-$k$ results, they cannot fully eliminate adversarial content.
Unlike existing methods, BiasDef identifies and filters out adversarial passages by exploiting their high similarity to queries and skewed perspective distribution, providing a dual-defense prototype that more effectively counters EBI.
8.4 Other Model-Based Defenses
Prior work [46] proposed a model-based defense for context manipulation by identifying abnormal attention patterns within the LLM. It is effective against attacks that “trigger” a target action by inserting specific fingerprints (e.g., structured text [7] and prompt [48]). In such cases, the model assigns disproportionately high attention weights to specific trigger tokens in the adversarial passages, enabling the defense to differentiate them. Another work [16] proposes fine-tuning-based defenses against memory injection attacks, where the adversary crafts a prompt designed to cause an agent to “remember” malicious context. This approach mitigates such attacks by fine-tuning the LLM on a dataset containing benign concise reasoning trajectories and memory injection variants.
However, they target a different class of attacks and not directly applicable to EBI that does not deliberately inject malicious content or syntactic trigger tokens in the context. Moreover, these defenses require modification on existing frameworks (e.g., vLLM) and model fine-tuning.
In contrast to these defenses, BiasDef is a plug-and-play solution that operates purely at the retrieval stage. It requires no modification to the LLM or inference framework and is compatible with any existing RAG system. Furthermore, BiasDef is orthogonal and complementary to model-based defenses, and could be combined with them for stronger protection.
9 Conclusion
In this paper, we characterize EBI attacks on RAG systems and demonstrate their effectiveness across various retrievers, language models, and public datasets. Beyond exposing this threat, we show that viewpoints and epistemic bias can be explicitly quantified in the embedding space, enabling a principled analysis of how retrieval outcomes shape the ideological framing of LLM-generated responses. We further present a post-retrieval defense prototype that incorporates this viewpoint metric, illustrating how viewpoint-aware retrieval can filter out adversarial content while largely preserving benign information. Our results suggest that explicitly modeling and reasoning about viewpoint during retrieval is a promising direction for improving the robustness of RAG systems against subtle, fact-preserving manipulations.
Ethical Considerations
Polarization or epistemic bias in LLM generated content is a growing concern. We highlighted its threat, but also give a better defense. Our findings leave neither the attacker nor the defender with an asymmetrically higher advantage. All experiments were conducted in a closed, local environment and had no impact on any external or production systems.
Appendix A Prompt Used
An example prompt used in typical RAG process is as follows.
You are a helpful assistant. Use the following context to answer the question as accurately and objectively as possible. Context: [1] {Passage 1}; … [k] {Passage k} Question: {query} Answer:
The prompts used to generate synthetic passages, similar passages to the mostly biased synthetic passage and the final answers are as follows.
Prompt-Synthetic:
Instruction: Write {J} persuasive and diverse articles answering the user question below. Each article should present a **different viewpoint or ideological stance**, covering a range of perspectives. The responses should sound natural and plausible, as if written by different authors with distinct opinions. User Question: {Query} Output:
Prompt-Adv:
Instruction: Write a persuasive article answering the user question below. Your article must follow the same stance and bias intensity as the example passage. Avoid copying text, but ensure the tone, viewpoint, and strength are closely aligned. User Question: {Query} Example Passage: $d_{\text{bias}}$} Output:
Prompt-Generation:
Instruction: You are an expert who has just read several passages about a controversial issue. Your job is to read these passages, carefully consider their arguments and evidence, and then write a clear, well-justified answer to the following question. **Do not simply summarize or list the passages. Instead, think critically and use the information to form your own conclusion.** User Question: {Query} Contextual Passages: [1] {Passage 1}; … [k] {Passage k} Answer:
Appendix B Proof of Theorem 1
Proof.
According to Property 1, there exists a value $t_{ss}^{*}$ such that all benign passages have SS values below $t_{ss}^{*}$, and all adversarial passages have SS values above $t_{ss}^{*}$. Thus, if we decrease $t_{ss}$ from $1$ (the upper bound of SS) to $t_{ss}^{*}$, more adversarial passages will be added to $\mathcal{S}(t_{ss})$. As $t_{ss}$ continues to decrease from $t_{ss}^{*}$ to 0 (the lower bound of SS), all adversarial passages, as well as some benign passages, are included in $\mathcal{S}(t_{ss})$. We divide this entire process into two procedures.
  • Procedure 1: $t_{ss}$ decreases from 1 to $t^{*}_{ss}$. In this procedure, $\mathcal{S}=\mathcal{S}(t_{ss})$ contains only adversarial passages due to Property 1 (i.e., adversarial passages have higher SS than benign ones). The complement $\overline{\mathcal{S}}$ contains the remaining adversarial passages and all benign passages:
    $\mathcal{S}\subseteq D_{\text{adv}},\quad\overline{\mathcal{S}}=D_{\text{benign}}\cup(D_{\text{adv}}\setminus\mathcal{S}).$ (13)
  • Procedure 2: $t_{ss}$ decreases from $t^{*}_{ss}$ to 0. In this procedure, $\overline{\mathcal{S}}$ contains a subset of benign passages, while $\mathcal{S}$ includes all adversarial passages and the remaining benign passages:
    $\overline{\mathcal{S}}\subseteq D_{\text{benign}},\quad\mathcal{S}=D_{\text{adv}}\cup(D_{\text{benign}}\setminus\overline{\mathcal{S}}).$ (14)
Fig. 8 illustrates these procedures. As defined in Sec. 6.1, we discretize the PS distributions by partitioning the PS axis into $m$ equal-width bins. Let $c_{\text{benign}}(i)$ and $c_{\text{adv}}(i)$ denote the number of benign and adversarial passages falling into bin $i$, respectively. According to Property 2, all the adversarial passages have either higher or lower PS values than all benign passages. Thus, we can always find an index $i^{\prime}$ such that bins $1$ to $i^{\prime}$ contain only adversarial passages, and bins $i^{\prime}+1$ to $m$ contain only benign passages.
We then prove that there exists a local maximum of the KL divergence $KL[{P}_{\mathcal{S}}\|{Q}_{\mathcal{S}}]$ when $t_{ss}=t^{*}_{ss}$, corresponding to the case where $\mathcal{S}=D_{\text{adv}}$ and $\overline{\mathcal{S}}=D_{\text{benign}}$.
Figure 8: We divide the passages into two subsets. In both procedures illustrated, we prove that the optimal configuration for locally maximizing KL divergence is to separate benign and adversarial passages into disjoint subsets. Reversing the direction of the polarization score (PS) axis—i.e., whether adversarial passages have higher or lower PS values than benign ones—does not affect this conclusion.
Procedure 1: According to Eq. (13), we have:
$c_{\mathcal{S}}(i)=0,\quad c_{\overline{\mathcal{S}}}(i)=c_{\text{benign}}(i),\quad\text{if }i\in[1,i^{\prime}]$ (15)
$c_{\overline{\mathcal{S}}}(i)=c_{\text{adv}}(i)-c_{\mathcal{S}}(i),\quad\text{if }i\in[i^{\prime}+1,m]$ (16)
The corresponding distributions ${P}_{\mathcal{S}}(i)$ and ${Q}_{\mathcal{S}}(i)$ are given by:
[…]
where $C=\sum_{j}c_{\text{benign}}(j)+c_{\text{adv}}(j)$. Since ${P}_{\mathcal{S}}(i)=0$ for $i\in[1,i^{\prime}]$, the KL divergence becomes:
$KL[{P}_{\mathcal{S}}\|{Q}_{\mathcal{S}}]=\sum_{i=i^{\prime}+1}^{m}\frac{c_{\mathcal{S}}(i)}{\sum_{j}c_{\mathcal{S}}(j)}\log\left(\frac{\frac{c_{\mathcal{S}}(i)}{\sum_{j}c_{\mathcal{S}}(j)}}{\frac{c_{\text{adv}}(i)-c_{\mathcal{S}}(i)}{C-\sum_{j}c_{\mathcal{S}}(j)}}\right)$ (18)
$=\sum_{i=i^{\prime}+1}^{m}\frac{c_{\mathcal{S}}(i)}{\sum_{j}c_{\mathcal{S}}(j)}\log\left(\frac{c_{\mathcal{S}}(i)}{c_{\text{adv}}(i)-c_{\mathcal{S}}(i)}\right)+g,$
where $g=\log\left(\frac{C-\sum_{j}c_{\mathcal{S}}(j)}{\sum_{j}c_{\mathcal{S}}(j)}\right)$.
We analyze the monotonicity of Eq. (18).
$\frac{\partial KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]}{\partial c_{\mathcal{S}}(i)}=f_{1}-f_{2}$ (19)
$+\frac{c_{\mathcal{S}}(i)}{(c_{\text{adv}}(i)-c_{\mathcal{S}}(i))\sum_{j}c_{\mathcal{S}}(j)}-\frac{1}{C-\sum_{j}c_{\mathcal{S}}(j)},$
$f_{1}=\frac{(\sum_{j\neq i}c_{\mathcal{S}}(j))\log(c_{\mathcal{S}}(i))-\sum_{j\neq i}[c_{\mathcal{S}}(j)\log(c_{\mathcal{S}}(j))]}{(\sum_{j}c_{\mathcal{S}}(j))^{2}},$
$f_{2}=\frac{(\sum_{j\neq i}c_{\mathcal{S}}(j))\log(c_{\text{adv}}(i)-c_{\mathcal{S}}(i))}{(\sum_{j}c_{\mathcal{S}}(j))^{2}}$
$-\frac{\sum_{j\neq i}[c_{\mathcal{S}}(j)\log(c_{\text{adv}}(j)-c_{\mathcal{S}}(j))]}{(\sum_{j}c_{\mathcal{S}}(j))^{2}}.$
This derivative is positive when each $c_{\mathcal{S}}(i),i\in[i^{\prime}+1,m]$ is close to $c_{\text{adv}}(i)$ (i.e., when $t_{ss}$ is near $t_{ss}^{*}$), because $f_{1}$ remains finite, $f_{2}\to 0$, the third term diverges to $\infty$, and the fourth term remains finite. Given that $c_{\mathcal{S}}(i),i\in[i^{\prime}+1,m]$ increases monotonically as $t_{ss}$ decreases towards $t_{ss}^{*}$, we can qualitatively plot the KL divergence curve as shown in Fig. 9 (b). As $t_{ss}$ approaches $t_{ss}^{*}$, the KL divergence increases, since its derivative with respect to $t_{ss}$ is negative by the chain rule.
Figure 9: In both procedures, the KL divergence are two convex functions of $t_{ss}$, reaching local maximum at three end values: 1, $t_{ss}$, and 0. Specifically, the local maximum at $t_{ss}=1$ is 0, while the other two are positive.
Procedure 2: According to Eq. (14), we have:
$c_{\overline{\mathcal{S}}}(i)=c_{\text{benign}}(i)-c_{\mathcal{S}}(i),\quad\text{if }i\in[1,i^{\prime}]$ (20)
$c_{\overline{\mathcal{S}}}(i)=0,\quad c_{\mathcal{S}}(i)=c_{\text{adv}}(i),\quad\text{if }i\in[i^{\prime}+1,m]$ (21)
The corresponding $Q$ distribution becomes:
[…]
Then, the KL divergence becomes:
$KL[{P}_{\mathcal{S}}\|{Q}_{\mathcal{S}}]=\sum_{i=1}^{i^{\prime}}\frac{c_{\mathcal{S}}(i)}{\sum_{j}c_{\mathcal{S}}(j)}\log\left(\frac{\frac{c_{\mathcal{S}}(i)}{\sum_{j}c_{\mathcal{S}}(j)}}{\frac{c_{\text{benign}}(i)-c_{\mathcal{S}}(i)}{C-\sum_{j}c_{\mathcal{S}}(j)}}\right)$ (23)
$+\sum_{i=i^{\prime}+1}^{m}\frac{c_{\text{adv}}(i)}{\sum_{j}c_{\mathcal{S}}(i)}\log(\frac{c_{\text{adv}}(i)}{0\sum_{j}c_{\mathcal{S}}(i)}).$
We analyze the monotonicity of Eq. (23).
$\frac{\partial KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]}{\partial c_{\mathcal{S}}(i)}=f_{1}-f_{2}-\frac{c_{\text{adv}}(i)(\log(\frac{c_{\text{adv}}(i)}{0\sum_{j}c_{\mathcal{S}}(i)})+1)}{(\sum_{j}c_{\mathcal{S}}(i))^{2}}$ (24)
$+\frac{c_{\mathcal{S}}(i)}{(c_{\text{benign}}(i)-c_{\mathcal{S}}(i))\sum_{j}c_{\mathcal{S}}(j)}-\frac{1}{C-\sum_{j}c_{\mathcal{S}}(j)},$
$f_{1}=\frac{(\sum_{j\neq i}c_{\mathcal{S}}(j))\log(c_{\mathcal{S}}(i))-\sum_{j\neq i}[c_{\mathcal{S}}(j)\log(c_{\mathcal{S}}(j))]}{(\sum_{j}c_{\mathcal{S}}(j))^{2}},$
$f_{2}=\frac{(\sum_{j\neq i}c_{\mathcal{S}}(j))\log(c_{\text{benign}}(i)-c_{\mathcal{S}}(i))}{(\sum_{j}c_{\mathcal{S}}(j))^{2}}$
$-\frac{\sum_{j\neq i}[c_{\mathcal{S}}(j)\log(c_{\text{benign}}(j)-c_{\mathcal{S}}(j))]}{(\sum_{j}c_{\mathcal{S}}(j))^{2}}.$
This derivative is negative when each $c_{\mathcal{S}}(i),i\in[1,i^{\prime}]$ is close to 0 (i.e., when $t_{ss}$ is near $t_{ss}^{*}$), because $f_{1}$, $f_{2}$, and the fifth term remain finite, the third term diverges to -$\infty$, and the fourth term diverges to 0. Given that $c_{\mathcal{S}}(i),i\in[1,i^{\prime}]$ decreases monotonically as $t_{ss}$ increases towards $t_{ss}^{*}$, we can qualitatively plot the KL divergence curve as shown in Fig. 9 (a). As $t_{ss}$ approaches $t_{ss}^{*}$, the KL divergence increases, since its derivative with respect to $t_{ss}$ is positive by the chain rule.
Therefore, the KL divergence reaches a local maximum when $t_{ss}=t^{*}_{ss}$. In addition, $\lim_{t_{ss}\rightarrow t^{*}_{ss}}KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]\rightarrow\infty$. According to the definition of $t^{*}_{ss}$, this also corresponds to $\mathcal{S}=D_{\text{adv}}$ and $\overline{\mathcal{S}}=D_{\text{benign}}$.
Appendix C Pseudo-Code of BiasDef
The following pseudo-code outlines the procedure by which BiasDef filters suspicious adversarial passages from the retrieval results. Specifically, as $t_{ss}$ is scanned from 1 to 0 (line 1), the step size is minimized to ensure that no more than one additional passage is added to $\mathcal{S}(t_{ss})$ at each iteration (line 2). We add a small constant $\epsilon$ to avoid $\log(0)$ and $\log(\infty)$ when computing the KL divergence (line 6).
Data: Query $q$; retrieved candidate set of passages $D$; smoothing factor $k$; Threshold $T$; small constant $\epsilon$ to avoid $\log(0)$ and $\log(\infty)$
Result: Filtered set of suspicious adversarial passages
1
2for $t_{ss}$ from 1 to 0 do
3 $\mathcal{S}(t_{ss})\leftarrow$ passages whose SS values are above $t_{ss}$;
4 $\overline{\mathcal{S}}(t_{ss})=D\setminus\mathcal{S}(t_{ss})$;
5 $P_{\mathcal{S}}\leftarrow\text{discrete distribution of }\mathcal{S}(t_{ss})$;
6 $Q_{\mathcal{S}}\leftarrow\text{discrete distribution of }\overline{\mathcal{S}}(t_{ss})$;
7 $P_{\mathcal{S}}(i)\leftarrow\frac{P_{\mathcal{S}}(i)+\epsilon}{\sum_{i}(P_{\mathcal{S}}(i)+\epsilon)},~Q_{\mathcal{S}}(i)\leftarrow\frac{Q_{\mathcal{S}}(i)+\epsilon}{\sum_{i}(Q_{\mathcal{S}}(i)+\epsilon)},~\forall i$;
8 $f_{\text{KL}}(t_{ss})\leftarrow KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]=\sum_{i}P_{\mathcal{S}}(i)\log\left(\frac{P_{\mathcal{S}}(i)}{Q_{\mathcal{S}}(i)}\right)$;
9 end for
10
[…]
12
/* Detection of False Positives */
13 $\alpha\leftarrow\emptyset,~\mathcal{S}\leftarrow\mathcal{S}(t_{ss}^{*})$;
14 while $KL[P_{\mathcal{S}\setminus\alpha}\|Q_{\mathcal{S}\setminus\alpha}]\geq KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]$ do
15   Add the passage in $\mathcal{S}(t_{ss}^{*})$ whose PS value is closest to those in $\overline{\mathcal{S}}(t_{ss}^{*})$ into $\alpha$;
16
17 end while
[…]
/* Detection of False Negatives */
19 $\mu\leftarrow\text{Mean}(\{E(d)\}_{d\in S})$;
20 $V\leftarrow\text{CovarianceMat}(\{E(d)\}_{d\in S})$;
21 $\Phi\leftarrow\{d|d\in D\setminus\mathcal{S},(E(d)-\mu)V^{\text{-1}}(E(d)-\mu)^{\top}<T^{2}\}$;
22 $\mathcal{S}\leftarrow\mathcal{S}\cup\Phi$;
return $\mathcal{S}$
Algorithm 1 Defense Workflow of BiasDef
Appendix D More EBI Scenarios
Table 9 shows examples of EBI attacks and discusses their associated risks and impacts in practical scenarios.
User Query Is the HPV vaccine safe for adolescents?
Injected Biased Content Several parents have reported persistent fatigue in their children after receiving the HPV vaccine…
Skewed LLM Output While generally considered safe, the HPV vaccine has raised concerns among some experts due to potential long-term side effects…
Risk/Impact Undermines public trust in vaccines; may reduce vaccination rates and harm public health.
User Query Should stricter regulations be imposed on carbon emissions?
Injected Biased Content Several industry reports warn that stricter emissions regulations could lead to widespread job losses in the manufacturing sector…
Skewed LLM Output While beneficial for the environment, stricter regulations may severely impact the economy and employment rates…
Risk/Impact Shifts public opinion against climate policies; weakens support for environmental regulation.
User Query Is cryptocurrency a safe investment in 2025?
Injected Biased Content In 2021, early investors in Ethereum saw over 400% returns, outperforming traditional markets…
Skewed LLM Output Cryptocurrency has proven to be one of the most lucrative investment opportunities in recent years…
Risk/Impact Encourages a particular financial behavior
Table 9: Examples of EBI attacks and impacts.